OpenStudy (anonymous):

John goes to school and returns home at speeds x km/h and (x+1)km/h respectively. The school is 2km from John's home and the total time for the two journeys is 54 minutes. Which of the following equations can be used to find x? A. \(\frac{ x }{ 2 }+\frac{ x+1 }{ 2 }=\frac{ 54 }{ 60 }\) B. \(\frac{ 2 }{ x }+\frac{ 2 }{ x+1 }=\frac{ 54 }{ 60 }\) C. \(\frac{ \frac{ 1 }{ 2 }(x+(x+1)) }{ 4 }=\frac{ 54 }{ 60 }\) D. \(\frac{ 4 }{ \frac{ 1 }{ 2 } (x+(x+1))}=\frac{ 54 }{ 60 }\) E. 2x+2(x+1)=\(\frac{ 54 }{ 60 }\)

OpenStudy (mathstudent55):

\( \text{speed} = \dfrac{\text{distance}}{\text{time}} \) \( \text{time} = \dfrac{\text{distance}}{\text{speed}} \) \( \text{time (in hours) going to school} = \dfrac{2}{x} \) \( \text{time (in hours) returning from school} = \dfrac{2}{x + 1}\) \( \text{total travel time (in hours)} = \dfrac{54}{60} \) \( \dfrac{2}{x} + \dfrac{2}{x + 1} = \dfrac{54}{60} \)

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm